A population consists of three numbers (3,5,9).Consider all possible samples of size n=2 which can be drawn WITHOUT REPLACEMENT from the population.

Questions to be answered:

Population mean

Population variance

Population standard deviation

Mean of the samples and mean of the sampling distribution of the means


1
Expert's answer
2021-06-24T08:38:33-0400

Given population = 3,5,9


Population mean "\mu =\dfrac{3+5+9}{3}=\dfrac{17}{3}=5.66"


Population variance


"\sigma^2= \sum \dfrac{ (x-\mu)^2}{n}=\dfrac{(3-5.66)^2+(5-5.66)^2+(9-5.66)^2}{3}=\dfrac{18.6}{3}=6.2"


Population standard deviation "\sigma=\sqrt{ Variance}=\sqrt{6.2}=2.49"



The mean of the samples are shown in the given table-



Total samples=3


mean of the sampling distribution of the means"=\dfrac{ \sum \text{sample mean}}{\text{total sample}}=\dfrac{17}{3}=5.66"